# Linear Discriminant Regularized Regression

### v2

@@ 4.2 Low-rank scenarios
In this section we evaluate our proposed procedure when the rank of the conditional mean matrix $M$ is low and hence $B$ is also low-rank. We consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Appendix A for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank.
Specifically, let the within-class covariance matrix $[\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$. To generate $M$, for some scalar $\eta>0$ and some randomly generated orthogonal matrix $A\in\mathbb{R}^{p\times r}$ with $A^{\top}A=\mathbf{I}_{r}$, we set $M=\eta A\alpha$ with entries of $\alpha\in\mathbb{R}^{r\times L}$ generated i.i.d. from $N(0,32/r)$. The quantity $\eta$ controls the magnitude of separation between the conditional means.
We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$. For each setting, we fix $n=1000,L=10,r=3,p=100$ and $\eta=1$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 2 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
@@
Figure 2: The averaged misclassification errors of each algorithm.
@@ 5 Real data analysis
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT). These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma. Each sample has $p=2,308$ gene expressions as features. The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004). The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes. We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction. We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$.
For all four data sets, the features are centered to have zero mean. Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data. This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
@@
LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA
Ramaswamy 30.7(0.051) 14.4(0.054) 22.6(0.048) 40.0(0.058) 30.4(0.058) 16.7(0.05)
SRBCT 0.7(0.022) 0.3(0.011) 0.1(0.007) 4.4(0.052) 1.0(0.019) 0.0(0.0)
Lymphoma 0.3(0.012) 0.5(0.017) 0.4(0.015) 1.9(0.034) 1.3(0.025) 1.3(0.025)
Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014)
Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison. From Table 1, we see that our procedure LDRR-F is close to have the best performance across all datasets. Choice of the best penalty depends on the data structure. For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent. For the Football dataset with $p=118$ and $L=10$, the selected rank is found to be $\widehat{r}=3$. Furthermore, in Figure 3, we show that LDRR-F with $K=2$ in (2.16) can be used for visualization in the 2D space of the first two discriminant directions. The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
Figure 3: Visualization of Ramaswamy, SRBCT, Lymphoma and Football data points in the space of the first two discriminant vectors.
@@ Acknowledgements
Appendix A Additional simulation study when the conditional mean matrix is low-rank and $\Sigma_{w}$ is approximately low-rank
In this setting, we consider the setting in Remark 1 and also compare with the method (PCLDA) in Bing and Wegkamp (2023). As mentioned in Remark 1, PCLDA is a two-step procedure similar to LDRR-F but with the first step done in a unsupervised way by using PCA. We generate the data according to model (1.1) with $M=\eta A\alpha$ as in Section 4.2 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $[\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.1 with $\rho=0.2$ and its diagonal elements from Uniform(0,1). The quantity $\eta$ represents the signal-to-noise ratio (SNR) between the low-dimensional signal $\eta AZ$ and the noise $W$. As pointed out in Bing and Wegkamp (2023), classification becomes easier as $\eta$ increases.
We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$. For all settings, we fix $n=1000,L=10,r=3,p=100$ and $\eta=2$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 4 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
Figure 4: The averaged misclassification errors of each algorithm.
Appendix B Proofs
Throughout the proofs, we will use the following notation. Define
$\Delta={M}^{\top}\Sigma_{w}^{-1}{M}$
and write
$\Delta_{\infty}=\max_{\ell\in[L]}\Delta_{\ell\ell},\qquad\Delta_{\rm op}=\|\Delta\|_{\rm op}.$
Further write
@@ Appendix B Proofs
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (B.1)
for some absolute constant $C>0$. Lemma 21 in Appendix D ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (B.2)
@@
B.1 Proof of the equivalence between (1.2) and (2.14)
Proof.
Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
so that, for any $\ell\in[L]$,
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
which, in view of (1.2) and (2.1), completes the proof. ∎
B.2 Proof of 2
For any $x\in\mathbb{R}^{p}$, write
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
@@ B.2 Proof of 2
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (B.3)
@@
We condition on ${\bm{D}}$ throughout the proof. The law of $\mathbb{E}$ and $\mathbb{P}$ is with respect to the randomness of $(X,Y)$, which is independent of ${\bm{D}}$. By definition, the excess risk $\mathcal{R}(\widehat{g})$ equals to
@@
Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
@@
[…]
@@
$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$
@@
For the first term, we have
$\sum_{k\neq\ell}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$ $\leq(L-1)\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=(L-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$ (B.4)
@@
${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (B.5)
@@
By the mean-value theorem, the quantity in (B.5) is no greater than
${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})^{2}$ $\leq{1\over 2}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
Together with (B.4) and $\pi_{k}\leq C/L$, the proof is complete. ∎
B.3 Proof of Theorem 3
@@
$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$ (B.6)
Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (B.1) which implies (B.2), an upper bound for the last term on the right of (B.6). Recall that $\mathbb{P}\{\mathcal{E}_{\pi}\}\geq 1-2n^{-2}$ and note that we consider both the randomness of $(X,Y)$ and that of ${\bm{D}}$ in this proof.
For the first term in (B.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and take a union bound over $\ell\in[L]$ to conclude that the event
$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~{}\text{for all }\ell\in[L]\right\}$ (B.7)
has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (B.6)
@@
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\ell\ell}}$ (B.8)
@@
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (B.9)
holds with probability at least $1-2Ln^{-2}$. The last step uses (B.8). Collecting (B.2), (B.7), (B.8) and (B.3) completes the proof. ∎
B.4 Proof of 4

### v3

@@ 4.3 Low-rank scenarios
In this section we evaluate our proposed procedure when the rank of the conditional mean matrix $M$ is low and hence $B$ is also low-rank.
4.3.1 Low-rank model (1)
We first consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Section 4.3.2 for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank. Specifically, let the within-class covariance matrix $\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$. To generate $M$, for some scalar $\eta>0$ and some randomly generated orthogonal matrix $A\in\mathbb{R}^{p\times r}$ with $A^{\top}A=\mathbf{I}_{r}$, we set $M=\eta A\alpha$ with entries of $\alpha\in\mathbb{R}^{r\times L}$ generated i.i.d. from $N(0,32/r)$. The quantity $\eta$ controls the magnitude of separation between the conditional means.
We examine how the performance of all algorithms depends on various parameters by varying one at a time: the number of classes $L\in\{5,15,25,35,45\}$; the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the separation scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$. For each setting, we fix $n=1000,L=10,r=3,p=100$ and $\eta=1$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 2 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
@@ 4.3.1 Low-rank model (1)
Figure 2: The averaged misclassification errors in low-rank model (1).
4.3.2 Low-rank model (2).
In this setting, we consider the setting in Remark 1 and also compare with the method (PCLDA) in Bing and Wegkamp (2023). As mentioned in Remark 1, PCLDA is a two-step procedure similar to LDRR-F but with the first step done in a unsupervised way by using PCA. We generate the data according to model (3.1) with $M=\eta A\alpha$ as in Section 4.3 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.2 with $\rho=0.2$ and its diagonal elements from Uniform(0,1). The quantity $\eta$ represents the signal-to-noise ratio (SNR) between the low-dimensional signal $\eta AZ$ and the noise $W$. As pointed out in Bing and Wegkamp (2023), classification becomes easier as $\eta$ increases.
We examine how the performance of all algorithms depends on different parameters by varying the parameters one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$. For all settings, we fix $n=1000,L=10,r=3,p=100$ and $\eta=2$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 3 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
Figure 3: The averaged misclassification errors in low-rank model (2).
@@ 5 Real data analysis
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The second dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004). The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes. We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction. We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$. Brain A is another dataset used in (Dettling, 2004). It has sample size $n=42$ with $p=5597$ features and contains $L=5$ different tumor types. For all four data sets, the features are centered to have zero mean. Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data. This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
@@
LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA SLDTR MSDA
Ramaswamy 22.3(0.051) 15.9(0.042) 22.9(0.044) 37.4(0.052) 28.6(0.043) 17.1(0.038) 17.9(0.045) 30.4(0.067)
Lymphoma 1.9(0.029) 3.1(0.051) 1.9(0.029) 3.1(0.044) 2.0(0.029) 1.9(0.029) 1.8(0.037) 5.9(0.050)
Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014) 34.11(0.011) 31.6 (0.011)
Brain A 24.3(0.106) 16.3(0.105) 13.8(0.082) 18.8 (0.107) 18.8(0.090) 19.3(0.080) 12.8(0.062) 39.0(0.134)
Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison. From Table 1, we observe that our procedure LDRR-F exhibits relatively better performance across all datasets. Choice of the best penalty depends on the data structure. For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent (the selected rank is found to be $\widehat{r}=3$ in the Football dataset with $p=118$ and $L=10$). We remark that in practice one can also use cross-validation to find the most suitable choice of penalty. Furthermore, in Figure 4, we show that LDRR-F with $K=2$ in (4.3) can be used for visualization in the 2D space of the first two discriminant directions. The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
Figure 4: Visualization in the space of the first two discriminant vectors.
@@ Acknowledgements
Appendix A contains the main proofs. Technical lemmas are stated in Appendix B.
Appendix A Main proofs
Throughout the proofs, we will use the following notation. Write
$\Delta_{\infty}=\max_{\ell\in[L]}{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell},\qquad\Delta_{\rm op}=\|{M}^{\top}\Sigma_{w}^{-1}{M}\|_{\rm op}.$
For future reference, we note that (under $\mathbb{E}(X)=0_{p}$)
${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})$
so that $\Delta_{\infty}\asymp\Delta$ under 2. Further, we write
@@ Appendix A Main proofs
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (A.1)
for some absolute constant $C>0$. Lemma 21 in Section B.2 ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (A.2)
@@
A.1 Proof of 2
Recall that, for any $x\in\mathbb{R}^{p}$,
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g^{*}(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
@@ A.1 Proof of 2
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (A.3)
@@
[…]
@@
Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y=k$ for each $k\in[L]$. We find that
@@
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\int_{\widehat{g}=\ell,g^{*}=k}\pi_{k}f_{k}(x)\left(1-\exp\left\{G^{(\ell|k)}(x)\right\}\right)dx$
@@
[…]
@@
For the first term, it is bounded from above by
$\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=\mathbb{P}(\mathcal{E}_{t}^{c}).$ (A.4)
@@
${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (A.5)
@@
By the mean-value theorem and 2, the quantity in (A.5) is no greater than
${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}{(t_{k}+t_{\ell})^{2}\over\sqrt{c\Delta}}$ $\leq{1\over 2\sqrt{c\Delta}}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2\sqrt{c\Delta}}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
Together with (A.4) and $\pi_{k}\leq C/L$, the proof of (3.6) is complete.
The bound in (3.7) follows by noting that
${1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{E}\left[\mathbbm{1}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y=k\right\}\right]$
$\leq{1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{P}\left\{\widehat{g}(X)=\ell\mid Y=k\right\}$
$\leq\max_{\ell\in[L]}t_{\ell}.$
The proof is complete. ∎
A.2 Proof of Theorem 3
@@
[…]
Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (A.1) which implies (A.2), an upper bound for the last term on the right of (A.6). Recall that $\mathbb{P}\{\mathcal{E}_{\pi}\}\geq 1-2n^{-2}$ and note that we consider both the randomness of $(X,Y)$ and that of $\bm{D}$} in this proof.
For the first term in (A.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}\leq\Delta_{\infty}$ and take a union bound over $\ell\in[L]$ to conclude that the event
$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~\text{for all }\ell\in[L]\right\}$ (A.7)
has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (A.6)
@@
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}}$ (A.8)
@@
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (A.9)
holds with probability at least $1-2Ln^{-2}$. The last step uses (A.8). Collecting 2, (A.2), (A.7), (A.8) and (A.2) and $\Delta_{\infty}\lesssim\Delta$ completes the proof. ∎
A.3 Proof of 4